# LUCI in the surface code with dropouts

### v1

@@ Abstract
Recently, usage of detecting regions facilitated the discovery of new circuits for fault-tolerantly implementing the surface code. Building on these ideas, we present LUCI, a framework for constructing fault-tolerant circuits flexible enough to construct aperiodic and anisotropic circuits, making it a clear step towards quantum error correction beyond static codes. We show that LUCI can be used to adapt surface code circuits to lattices with imperfect qubit and coupler yield, a key challenge for fault-tolerant quantum computers using solid-state architectures. These circuits preserve spacelike distance for isolated broken couplers or isolated broken measure qubits in exchange for halving timelike distance, substantially reducing the penalty for dropout compared to the state of the art and creating opportunities in device architecture design. For qubit and coupler dropout rates of $1\%$ and a patch diameter of 15, LUCI achieves an average spacelike distance of 13.1, compared to 9.1 for the best method in the literature. For a SI1000(0.001) circuit noise model, this translates to a 36×36$\times$36 × improvement in median logical error rate per round, a factor which increases with device performance. At these dropout and error rates, LUCI requires roughly $25\%$ fewer physical qubits to reach algorithmically relevant one-in-a-trillion logical codeblock error rates.
I Introduction
In order to reach the error rates necessary for large-scale quantum algorithms, we will require the use of quantum error correction (QEC). By most estimates, such a computer would require thousands of logical qubits, each composed of hundreds to thousands of physical qubits von Burg et al. (2021); Lee et al. (2021); Kim et al. (2022); Shokrian Zini et al. (2023); Rubin et al. (2023). In solid-state architectures, fabrication errors can lead to a number of failure modes that result in broken qubits, as well as broken couplers in architectures that use them, like Google’s Sycamore and USTC’s Zuchongzhi,  Google Quantum AI (2019); Wu et al. (2021); Zhu et al. (2022).
@@ I Introduction
In addition, transient errors can significantly degrade qubit or coupler performance for long time periods, such as drifting Two-Level Systems (TLSs) Klimov et al. (2018). We refer to qubits or couplers that are either permanently or temporarily unusable in the context of a QEC protocol as dropouts. In this manuscript we will focus on a simplified failure model, where qubits and couplers are missing from the grid with independent probabilities $p_{q}$ and $p_{c}$, respectively. In Fig. 1 we show an example dropout grid for a chip designed to fit a distance-9 surface code, where around 2% of the qubits and couplers are missing.
@@
Figure 2: Detection region slices in the Surface Code. Adapted from McEwen et al. (2023). (top) Zoomed in view of the circuit used to measure a standard surface code, with angled CNOT gates used to indicate gates which connect to qubits not labeled. (bottom) Timelike slices of the detecting regions of the surface code circuit at the indicated point. Crucially, the mid-cycle state of surface code circuit is an unrotated surface code. In this figure and throughout this manuscript, blue(red) will be used to indicate Z(X)-type Pauli operators unless specified otherwise.
@@
There are a number of methods known for adapting quantum error correction protocols to grids with qubit and coupler dropout Auger et al. (2017); Strikis et al. (2023); Siegel et al. (2023); Lin et al. (2024); Wei et al. (2024). In most cases, isolated broken data qubits reduce distance by one in both directions, isolated broken measure qubits reduce distance by two. Isolated broken couplers are fixed by removing the data qubit which interacts via that coupler. More substantial differences between the methods arise when dealing with multiple broken components in a small area, with the method described in Auger et al. (2017), and later rediscovered by Wei et al. (2024), better preserving the functional parts of the chip.
@@
In this manuscript we will discuss a new method for building quantum error correction circuits for grids with dropouts, based on some of the ideas first introduced in McEwen et al. (2023). These circuits are constructed from a small set of different rounds, each starting and ending in a modified mid-cycle state of the surface code. We introduce a visual language for indicating which mid-cycle stabilizers are measured in that round, and via which operations. Individual mid-cycle stabilizer measurements are described by shapes, including L-, U-, C-, and I-shapes, amongst others, so these diagrams are referred to as LUCI diagrams.
@@
Figure 3: LUCI Diagrams and their circuit interpretation. (left) LUCI diagram showing the two rounds of a distance-5 three-coupler surface code circuit McEwen et al. (2023). Colors indicate the mid-cycle stabilizers of the surface code being measured, with X stabilizers in red and Z stabilizers in blue. The gray shapes inside the squares indicate how that stabilizer will be measured. (right) Circuit compilations for U-shapes on X-type (red) and Z-type (blue) squares. The colored regions depict the detecting region contracting and being measured out, and then a new detecting region expanding from a reset. Other shapes, like the L’s on the top and left boundaries, are formed by removing the appropriate CNOT gates from these U-shape compilations.
@@ II LUCI
II.1 Detecting regions and the mid-cycle state
@@
In Ref. McEwen et al. (2023), the authors introduce an approach to error correction circuits where the standard stabilizers of the quantum error correcting code are propagated through the circuit to form so-called “detecting regions”. Detecting regions capture the extent of the stabilizer in spacetime, where Pauli errors are detectable by looking at a specific set of measurement outcomes. By approaching the problem of QEC circuit construction as the task of covering the circuit with appropriate detecting regions (as opposed to constructing the circuit directly from a code), we reveal additional freedom to manipulate the circuit and improve performance.
@@
Cross-sections of the detecting regions at key points during a a standard surface code circuit are shown in Fig. 2. In the bulk rounds of the circuit detecting regions survive for two rounds, starting at measure qubit initialization, expanding into the typical code stabilizer in the first round of entangling gates, and then contracting back to be terminated in the second round. Detectors can be formed by combining all the measurements that a given detecting region terminates on, which for the standard circuit is simply a comparison of subsequent measurements on the relevant measure qubit. In LUCI circuits, we produce more complex structures of detecting regions that span additional rounds and terminate on spatially separate measurements. As such, the detecting region picture is particularly important in understanding and building the detectors for our circuits, which is described in Sec. II.3.
@@
One insight of Ref. McEwen et al. (2023) is that the mid-cycle state of the standard surface code state is an unrotated surface code state on both measure and data qubits. This allowed the construction of novel surface code circuits by measuring the mid-cycle stabilizers in different ways before returning to the same state. In this view, the circuit is constructed from mid-cycle state to mid-cycle state, with half-rounds at the beginning and end of the circuit to get to the usual initial and final states of a surface code circuit. LUCI approaches the task of constructing circuits around dropouts in a similar way, using the mid-cycle surface code state as a “home base” to return to.
@@
The LUCI construction in particular has similarities to the three-coupler surface code circuit from Ref. McEwen et al. (2023). Using a pair of CNOT gates, a weight-4 bulk mid-cycle stabilizer is “folded” onto the pair of qubits along one edge of the square. This weight-2 operator is then folded again using a single CNOT gate, after which it is measured and reset, before being unfolded back into the original weight-4 footprint. Adjacent mid-cycle stabilizers can be measured simultaneously using a “snake” configuration, where the gates are positioned such that the CNOT gates for the initial circuit layer fold both stabilizers simultaneously. This snake configuration will be seen again in Sec. II.2, as it will be the rule we use to fit LUCI shapes together correctly.
II.2 LUCI diagrams
@@
LUCI diagrams provide a visual language for describing circuit constructions starting at the mid-cycle state. In order to construct a LUCI circuit, we start by finding a set of mid-cycle stabilizer generators which are compatible with the dropout grid in question. Some of these stabilizers may be composed of multiple mid-cycle gauge operators multiplied together, as in other subsystem code dropout constructions. We then specify a set of shapes that describe how to measure these mid-cycle gauge operators and stabilizers over multiple rounds. Each of the rounds from mid-cycle to mid-cycle is described by a board in the LUCI diagram, an example of which is shown in Fig. 3. A LUCI circuit iterates through these boards sequentially, returning to the mid-cycle state as a reliable interface between rounds. Because the mid-cycle state is stable, the process of finding logical operators simply requires finding an error string in the mid-cycle state from one corner to another, and then propagating it through the circuit. The detector cross-section at measurement varies for each round, as illustrated in Appendix D, indicating that LUCI circuits implement a dynamic code.
@@
Figure 4: Compatibility of neighbouring LUCI shapes. Examples of adjacent squares which are compatible (top row) and incompatible (bottom two rows). Like for the LUCI diagrams, qubits in the circuit layouts are at the vertices of the square grid, with gate layers indicated by the numbers to match with the compilation in Fig. 3. The dial icon indicates a MRX or MRZ gate for red or blue squares, respectively. For the two incompatible examples, the gates highlighted in red show the collision which makes the diagram invalid.
An interesting note is that in a LUCI diagram, reversing the orientation of the second layer of CNOTs allows you to change switch the qubits that are measured without impacting compatibility. As a result, it is straightforward to build LUCI circuits which switch data qubit and measure qubit roles in successive rounds, which can be impactful for leakage errors McEwen et al. (2023); Camps et al. (2024). One could do this within a usual four-round cycle, or double the cycle to eight round and switch qubit roles for the second half of the set of eight. This flexibility is one of the primary advantages of LUCI, making it easy to modify a circuit to measure particularly leaky qubits more often than others without much difficulty, or whatever else suits the experimental conditions.
II.3 Generating Circuits from LUCI Diagrams
One method to guarantee all squares are measured when modifying diagrams is to use a four-coloring of the underlying square lattice of the mid-cycle state, shown in the seventh step of Fig. 8. Since the squares of a given color do not share any qubits, they can all be filled independently without any compatibility concerns, in this way LUCI allows us to measure shapes which would otherwise be incompatible by interleaving in time. By requiring that each square is filled in its highlighted round, we guarantee a fully measured mid-cycle state. Additional squares can also be measured in each round to minimize logical error rate, and in Sec. II.4 we build around a base diagram, to both maximize the number of measurements and make the circuits more structured, reducing calibration overhead.
@@ II.3 Generating Circuits from LUCI Diagrams
We show the structure for a representative detecting region in Fig. 5, For LUCI, we have to be more careful than in the standard surface code case. Unlike in the usual circuit, boards avoiding dropouts may pull detecting regions outwards before returning them into their usual position, leading to detectors that consist of a number of measurements on different qubits, as can be seen in other middle-out constructions McEwen et al. (2023); Shaw and Terhal (2024). A detecting region starts at a reset, after which it takes two CNOT layers to expand into a mid-cycle stabilizer. The next round generally does not measure this same mid-cycle stabilizer, and in this case the detecting region may be manipulated by the measurement of nearby mid-cycle stabilizers. The region is returned to its original position, possibly by including measurements and resets on the way. It then is folded and measured fully, completing the region. The corresponding detector combines all the measurements which touch the region.
II.4 Building an example LUCI circuit
@@
Figure 7: Finding the mid-cycle state.                                                                                                                                                 1. Starting from the original dropout grid in Fig 1, convert each broken qubit into four broken couplers connected to that qubit, then remove any qubit which has two perpendicular broken couplers attached to it. LUCI handles broken couplers by passing information around the broken coupler, which is impossible if an additional qubit or coupler is broken. As a result, information is trapped on the qubit isolated by the broken components, so this otherwise functional qubit must be removed. A qubit is removed on the left due to a broken coupler being too close to a broken qubit, and one is removed from the top right for having two incident broken couplers.                                                                                                                                                 2. Remove all broken qubits from mid-cycle stabilizers they touch. This will lead to some non-commuting Pauli operators, which we will refer to as mid-cycle gauge operators. Like in a subsystem code, the gauge operators commute with all stabilizers, but may anti-commute with other gauge operators. All remaining broken couplers do not prevent us from measuring the usual stabilizers, so they are not treated differently at this stage.                                                                                                                                                 3. Form super-stabilizers by grouping mid-cycle gauge operators into products which commute with all other gauge operators. Extending stabilizers into detecting regions keeps commutation relations, so doing this grouping at the mid-cycle gives the same results as grouping at the end-cycle. This grouping problem admits multiple solutions, so we choose the groupings that minimize super-stabilizer size. In the figure, grouped mid-cycle gauge operators are indicated by the red and blue dotted regions. As explained in Auger et al. (2017), this process can be thought of as combining stabilizer generators of the same type that touch a given broken qubit, forming new stabilizers which are not supported on the qubit in question. Broken components that are clustered or near the boundaries may lead to gauge operators which cannot be paired, and must be removed or handled differently. This is discussed further in App. B.                                                                                                                                                 4. Finding mid-cycle logical operators. We start by identifying the corners of the code, qubits which are in a single stabilizer of each type. In the figure these qubits are indicated by dark gray circles. Bare logical X(Z) operators, shown in red(blue), can be found by connecting two corners on opposite X(Z) boundaries of the mid-cycle state with operators that commute with every mid-cycle stabilizer and mid-cycle gauge operator. In the detecting region picture, these logical operators are sheets in the 2+1-dimensional view of the circuit, so identifying the logical operator at the mid-cycle fully defines the operator at all other time slices. Over the course of the circuit the logical operator will move due to CNOT gates, but return to the same support for each mid-cycle state.
@@
The described method is simply one way to build a LUCI diagram for a dropout grid. There are a number of optimizations that can be made on top of this method to target different error models and hardware constraints. As an example, more measurements can be inserted by flipping parts of some rows to alleviate incompatibility issues. However, rows that face each other lead to unevenly sized detecting regions, as one basis will get stretched more than the other. This produces uneven detection event fractions, causing logical error rate biases and worsened performance when decoding. Additional optimizations include removing boundary qubits which are only every used by single-qubit mid-cycle stabilizers and as edge qubits in weight-4 stabilizers, as the weight-4 stabilizer could be converted into an L-shape without damaging the code. This would reduce the footprint of the code without a penalty on performance, depending on the error model. In this paper we will keep these extra qubits to avoid additional complexity.
@@
For intuition as to why this method outperforms the currently known state of the art, we point to two key features. Firstly, the holes that are cut around broken components are much smaller than in previous methods since they exist in the mid-cycle state. This means that we do not have as many issues with nearby holes merging together, and suffer less of a performance penalty from individual dropouts. The super-stabilizers that are formed also are oriented such that they do not have additional extent along logical operators for measure qubits. In App. C we explain how this asymmetry is purely geometric, given that the data and measure qubit roles are arbitrary in LUCI. This lack of fixed qubit roles also make it very easy to trade roles when needed. In the case that the missing qubit would have been measured, nearby qubits end up doing double-duty to still measure relevant Pauli operators, much like the Surface-13 construction in Ref. Tomita and Svore (2014). This still has a penalty on performance, but preserves spacelike distance.
@@
In addition, LUCI circuits can be generated by randomly sampling arbitrary applicable shapes in the subgrids of step 7. The subgrid framework guarantees that the detecting regions only span 4 rounds, so this method could be used to create aperiodic and anisotropic error correcting circuits which still maintain spacelike distance and perform reasonably close to the usual surface code. Such circuits may not be relevant for most platforms, but are interesting in terms of vastly expanding the space of circuits usable for error correction, and could be useful for random compiling Beale and Wallman (2023); Jain et al. (2023).
III Results

### v2

@@ Abstract
Recently, usage of detecting regions facilitated the discovery of new circuits for fault-tolerantly implementing the surface code. Building on these ideas, we present LUCI, a framework for constructing fault-tolerant circuits flexible enough to construct aperiodic and anisotropic circuits, making it a clear step towards quantum error correction beyond static codes. We show that LUCI can be used to adapt surface code circuits to lattices with imperfect qubit and coupler yield, a key challenge for fault-tolerant quantum computers using solid-state architectures. These circuits preserve spacelike distance for isolated broken couplers or isolated broken measure qubits in exchange for halving timelike distance, substantially reducing the penalty for dropout compared to the state of the art and creating opportunities in device architecture design. For qubit and coupler dropout rates of $1\%$ and a patch diameter of 15, LUCI achieves an average spacelike distance of 13.1, compared to 9.1 for the best method in the literature. For a SI1000(0.001) circuit noise model, this translates to a $36\times$ improvement in median logical error rate per round, a factor which increases with device performance. At these dropout and error rates, LUCI requires roughly $25\%$ fewer physical qubits to reach algorithmically relevant one-in-a-trillion logical codeblock error rates.
1 Introduction
In order to reach the error rates necessary for large-scale quantum algorithms, we will require the use of quantum error correction (QEC). By most estimates, such a computer would require thousands of logical qubits, each composed of hundreds to thousands of physical qubits [1, 2, 3, 4, 5]. In solid-state architectures, fabrication errors can lead to a number of failure modes that result in broken qubits, as well as broken couplers in architectures that use them, like Google’s Sycamore and USTC’s Zuchongzhi [6, 7, 8].
In addition, transient errors can significantly degrade qubit or coupler performance for long time periods, such as drifting Two-Level Systems (TLSs) [9].
@@ 1 Introduction
We refer to qubits or couplers that are either permanently or temporarily unusable in the context of a QEC protocol as dropouts.
@@
Figure 2: Detection region slices in the Surface Code. Adapted from [10]. (top) Zoomed in view of the circuit used to measure a standard surface code, with angled CNOT gates used to indicate gates which connect to qubits not labeled. (bottom) Timelike slices of the detecting regions of the surface code circuit at the indicated point. Crucially, the mid-cycle state of surface code circuit is an unrotated surface code. In this figure and throughout this manuscript, blue(red) will be used to indicate Z(X)-type Pauli operators unless specified otherwise. For visual clarity, weight-1 operators are represented as circles slightly offset from the relevant qubit.
In this manuscript we will focus on a simplified failure model, where qubits and couplers are missing from the grid with independent probabilities $p_{q}$ and $p_{c}$, respectively. In Fig. 1 we show an example dropout grid for a chip designed to fit a distance-9 surface code, where around 2% of the qubits and couplers are missing.
@@
There are a number of methods known for adapting quantum error correction protocols to grids with qubit and coupler dropout [11, 12, 13, 14, 15, 16]. In most cases, isolated broken data qubits reduce distance by one in both directions, isolated broken measure qubits reduce distance by two. Isolated broken couplers are fixed by removing the data qubit which interacts via that coupler. More substantial differences between the methods arise when dealing with multiple broken components in a small area, with the method described in [11], and later rediscovered by [15], better preserving the functional parts of the chip.
@@
In this manuscript we will discuss a new method for building quantum error correction circuits for grids with dropouts, based on some of the ideas first introduced in [10]. These circuits are constructed from a small set of different rounds, each starting and ending in a modified mid-cycle state of the surface code. We introduce a visual language for indicating which mid-cycle stabilizers are measured in that round, and via which operations. Individual mid-cycle stabilizer measurements are described by shapes, including L-, U-, C-, and I-shapes, amongst others, so these diagrams are referred to as LUCI diagrams.
@@
Figure 3: LUCI Diagrams and their circuit interpretation. (left) LUCI diagram showing the two rounds of a distance-5 three-coupler surface code circuit [10]. Colors indicate the mid-cycle stabilizers of the surface code being measured, with X stabilizers in red and Z stabilizers in blue. The gray shapes inside the squares indicate how that stabilizer will be measured. (right) Circuit compilations for U-shapes on X-type (red) and Z-type (blue) squares. The colored regions depict the detecting region contracting and being measured out, and then a new detecting region expanding from a reset. Shades of gray in the shape denote which circuit layer the operation occurs in. Other shapes, like the L’s on the top and left boundaries, are formed by removing the appropriate CNOT gates from these U-shape compilations.
@@ 2 LUCI
2.1 Detecting regions and the mid-cycle state
@@
In Ref. [10], the authors introduce an approach to error correction circuits where the standard stabilizers of the quantum error correcting code are propagated through the circuit to form so-called “detecting regions". Detecting regions capture the extent of the stabilizer in spacetime, where Pauli errors are detectable by looking at a specific set of measurement outcomes. By approaching the problem of QEC circuit construction as the task of covering the circuit with appropriate detecting regions (as opposed to constructing the circuit directly from a code), we reveal additional freedom to manipulate the circuit and improve performance.
@@
Cross-sections of the detecting regions at key points during a standard surface code circuit are shown in Fig. 2. In the bulk rounds of the circuit detecting regions survive for two rounds, starting at measure qubit initialization, expanding into the typical code stabilizer in the first round of entangling gates, and then contracting back to be terminated in the second round. Detectors can be formed by combining all the measurements that a given detecting region terminates on, which for the standard circuit is simply a comparison of subsequent measurements on the relevant measure qubit. In LUCI circuits, we produce more complex structures of detecting regions that span additional rounds and terminate on spatially separate measurements. As such, the detecting region picture is particularly important in understanding and building the detectors for our circuits, which is described in Sec. 2.3.
@@
One insight of Ref. [10] is that the mid-cycle state of the standard surface code state is an unrotated surface code state on both measure and data qubits. This allowed the construction of novel surface code circuits by measuring the mid-cycle stabilizers in different ways before returning to the same state. In this view, the circuit is constructed from mid-cycle state to mid-cycle state, with half-rounds at the beginning and end of the circuit to get to the usual initial and final states of a surface code circuit. LUCI approaches the task of constructing circuits around dropouts in a similar way, using the mid-cycle surface code state as a “home base" to return to.
@@
The LUCI construction in particular has similarities to the three-coupler surface code circuit from Ref. [10]. Using a pair of CNOT gates, a weight-4 bulk mid-cycle stabilizer is “folded” onto the pair of qubits along one edge of the square. This weight-2 operator is then folded again using a single CNOT gate, after which it is measured and reset, before being unfolded back into the original weight-4 footprint. Adjacent mid-cycle stabilizers can be measured simultaneously using a “snake” configuration, where the gates are positioned such that the CNOT gates for the initial circuit layer fold both stabilizers simultaneously. This snake configuration will be seen again in Sec. 2.2, as it will be the rule we use to fit LUCI shapes together correctly.
2.2 LUCI diagrams
@@
LUCI diagrams provide a visual language for describing circuit constructions starting at the mid-cycle state. In order to construct a LUCI circuit, we start by finding a set of mid-cycle stabilizer generators which are compatible with the dropout grid in question. Some of these stabilizers may be composed of multiple mid-cycle gauge operators multiplied together, as in other subsystem code dropout constructions. We then specify a set of shapes that describe how to measure these mid-cycle gauge operators and stabilizers over multiple rounds. Each of the rounds from mid-cycle to mid-cycle is described by a board in the LUCI diagram, an example of which is shown in Fig. 3. A LUCI circuit iterates through these boards sequentially, returning to the mid-cycle state as a reliable interface between rounds. Because the mid-cycle state is stable, the process of finding logical operators simply requires finding an Pauli string in the mid-cycle state from one corner to another, and then propagating it through the circuit. The detector cross-section at measurement varies for each round, as illustrated in Appendix C, indicating that LUCI circuits implement a dynamic code.
@@
Figure 4: Compatibility of neighboring LUCI shapes. Examples of adjacent squares which are compatible (top row) and incompatible (bottom two rows). Like for the LUCI diagrams, qubits in the circuit layouts are at the vertices of the square grid, with gate layers indicated by the numbers to match with the compilation in Fig. 3. The dial icon indicates a MRX or MRZ gate for red or blue squares, respectively. For the two incompatible examples, the gates highlighted in red show the collision which makes the diagram invalid.
@@
An interesting note is that in a LUCI diagram, reversing the orientation of the second layer of CNOTs allows you to switch the qubits that are measured without impacting compatibility. As a result, it is straightforward to build LUCI circuits which switch data qubit and measure qubit roles in successive rounds, which can be impactful for leakage errors [10, 17]. One could do this within a usual four-round cycle, or double the cycle to eight round and switch qubit roles for the second half of the set of eight. This flexibility is one of the primary advantages of LUCI, making it easy to modify a circuit to measure particularly leaky qubits more often than others without much difficulty, or whatever else suits the experimental conditions.
2.3 Generating Circuits from LUCI Diagrams
One method to guarantee all squares are measured when modifying diagrams is to use a four-coloring of the underlying square lattice of the mid-cycle state, shown in the seventh step of Fig. 8. Since the squares of a given color do not share any qubits, they can all be filled independently without any compatibility concerns, in this way LUCI allows us to measure shapes which would otherwise be incompatible by interleaving in time. By requiring that each square is filled in its highlighted round, we guarantee a fully measured mid-cycle state. Additional squares can also be measured in each round to minimize logical error rate, and in Sec. 2.4 we build around a base diagram, to both maximize the number of measurements and make the circuits more structured, reducing calibration overhead. This method guarantees a valid LUCI diagram for arbitrary configurations of dropouts.
@@ 2.3 Generating Circuits from LUCI Diagrams
We show the structure for a representative detecting region in Fig. 5, For LUCI, we have to be more careful than in the standard surface code case. Unlike in the usual circuit, boards avoiding dropouts may pull detecting regions outwards before returning them into their usual position, leading to detectors that consist of a number of measurements on different qubits, as can be seen in other middle-out constructions [10, 18]. A detecting region starts at a reset, after which it takes two CNOT layers to expand into a mid-cycle stabilizer. The next round generally does not measure this same mid-cycle stabilizer, and in this case the detecting region may be manipulated by the measurement of nearby mid-cycle stabilizers. The region is returned to its original position, possibly by including measurements and resets on the way. It then is folded and measured fully, completing the region. The corresponding detector combines all the measurements which touch the region.
2.4 Building an example LUCI circuit
@@
Figure 7: Finding the mid-cycle state.                                   1. Starting from the original dropout grid in Fig 1, convert each broken qubit into four broken couplers connected to that qubit, then remove any qubit which has two perpendicular broken couplers attached to it. LUCI handles broken couplers by passing information around the broken coupler, which is impossible if an additional qubit or coupler along that path is broken. As a result, qubits isolated by multiple broken components in a single mid-cycle stabilizer are removed. A qubit is removed on the left due to a broken coupler being too close to a broken qubit, and one is removed from the top right for having two incident broken couplers.                                   2. Remove all broken qubits from mid-cycle stabilizers they touch. This will lead to some non-commuting Pauli operators, which we will refer to as mid-cycle gauge operators. Like in a subsystem code, the gauge operators commute with all stabilizers, but may anti-commute with other gauge operators. All remaining broken couplers do not prevent us from measuring the usual stabilizers, so they are not treated differently at this stage.                                   3. Form super-stabilizers by grouping mid-cycle gauge operators into products which commute with all other gauge operators. Extending stabilizers into detecting regions keeps commutation relations, so doing this grouping at the mid-cycle gives the same results as grouping at the end-cycle. This grouping problem admits multiple solutions, so we choose the groupings that minimize super-stabilizer size. In the figure, grouped mid-cycle gauge operators are indicated by the red and blue dotted regions. As explained in [11], this process can be thought of as combining stabilizer generators of the same type that touch a given broken qubit, forming new stabilizers which are not supported on the qubit in question. Broken components that are clustered or near the boundaries may lead to gauge operators which cannot be paired, and must be removed or handled differently. This is discussed further in App. B.                                   4. Finding mid-cycle logical operators. We start by identifying the corners of the code, qubits which are in a single stabilizer of each type. In the figure these qubits are indicated by dark gray circles. Bare logical X(Z) operators, shown in red(blue), can be found by connecting two corners on opposite X(Z) boundaries of the mid-cycle state with operators that commute with every mid-cycle stabilizer and mid-cycle gauge operator. In the detecting region picture, these logical operators are sheets in the 2+1-dimensional view of the circuit, so identifying the logical operator at the mid-cycle fully defines the operator at all other time slices. Over the course of the circuit the logical operator will move due to CNOT gates, but return to the same support for each mid-cycle state.
@@
The described method is simply one way to build a LUCI diagram for a dropout grid. There are a number of optimizations that can be made on top of this method to target different error models and hardware constraints. As an example, more measurements can be inserted by flipping the orientations of entire rows to alleviate incompatibility issues. However, shapes with are oriented opposite to each other lead to situations in which detecting regions of one type are stretched beyond the usual size, while the other basis is unstretched. This produces uneven detection event fractions, causing logical error rate biases and worsened performance when decoding. Additional optimizations include removing boundary qubits which are only ever used by single-qubit mid-cycle stabilizers and as edge qubits in weight-4 stabilizers, as the weight-4 stabilizer could be converted into an L-shape without damaging the code. This would reduce the footprint of the code without a penalty on performance, depending on the error model. In this paper we will keep these extra qubits to avoid additional complexity.
Figure 9: Explaining distance loss from qubit dropouts. The mid-cycle state for a grid with two qubit dropouts. The right side dropout creates Z(X)-type stabilizers which are extended vertically(horizontally). This means that the X(Z)-type logical operator, shown in red(blue), can use the extended stabilizer as a shortcut to reduce the logical distance by one. In contrast, the left side dropout has stabilizers extended parallel to the logical operator of their own type, so they do not reduce distance.
@@
For intuition as to why this method outperforms the currently known state of the art, we point to two key features. Firstly, the holes that are cut around broken components are much smaller than in previous methods since they exist in the mid-cycle state. This means that we do not have as many issues with nearby holes merging together, and suffer less of a performance penalty from individual dropouts. The super-stabilizers that are formed also are oriented such that they do not have additional extent along logical operators for measure qubits. In Fig. 9, we show the orientations of super-stabilizers in the mid-cycle state for a grid with two missing qubits. It can be seen that one qubit has super-stabilizers formed such that the X(Z)-type super-stabilizer is elongated in the direction of the Z(X)-type logical operator, reducing distance in both cases, while the other qubit has the super-stabilizers elongated perpendicular to the relevant logical operator. While we explain this behavior using the language of data and measure qubits for simplicity, the cause of this behavior is geometric, and independent on whether one compiles a circuit that would measure the missing qubit or not.
@@
In addition, LUCI circuits can be generated by randomly sampling arbitrary applicable shapes in the subgrids of step 7. The subgrid framework guarantees that the detecting regions only span 4 rounds, so this method could be used to create aperiodic and anisotropic error correcting circuits which still maintain spacelike distance and perform reasonably close to the usual surface code. Such circuits may not be relevant for most platforms, but are interesting in terms of vastly expanding the space of circuits usable for error correction, and could be useful for random compiling [19, 20].
3 Results